# LOLA: LLM-Assisted Online Learning Algorithm for Content Experiments

### v2

@@ D.2.2 DistilBERT LoRA Fine-Tuning
Figure A10: The classification pipeline using the DistilBERT model. Headlines are first tokenized, and each token is converted into an embedding vector. For illustrative purposes, only two Transformer layers are shown. The figure highlights the use of bidirectional attention in DistilBERT, whereby each token can incorporate information from any token. Connections between Transformer blocks represent the attention mechanism, and only a subset of these links are displayed for simplicity. The output from the first token (CLS token) is fed into a classification head, which computes the probability that one headline is more impactful than the other.
Another difference is that DistilBERT is a smaller model compared to Llama-3-8b. DistilBERT has around 66 million parameters, which is only 0.94% of the parameters of Llama-3-8b. Due to the smaller model size, we are able to conduct full parameter updates during the fine-tuning process without using LoRA. We use the same learning rate scheduling, number of training epochs, and train-test set split as in the experiment detailed in §$\S$3.3.2. The accuracy of DistilBERT LoRA fine-tuning on the same test dataset is around 75%. Thus, we find that the newer Llama-3-8b model offers better performance upon fine-tuning. We refer interested readers to our code repository for further details on the fine-tuning process.
@@ Appendix E Cost Comparison of Pure-LLM-Based Methods
Prompt \$21.24 (10,000 GPT prompts) 121.5 mins for 10,000 GPT prompts (82 prompt requests per min) Low
Embedding \$0.18 (text-embedding-3-large) 26 mins for embedding all 77,245 headlines (3,000 embedding requests per min) Moderate
@@
Table A2: Summary of cost, run-time, and effort for different methods. Note that the reported run-time for Prompt-based (GPT-3.5 and GPT-4) and embedding-based methods refer to the time needed to obtain responses or embeddings from OpenAI. Note that the prompt and embedding response time can vary based on the API usage tiers, with our results specifically derived from Tier 1.
For the prompt-based method utilizing OpenAI API, the cost was \$1.13 for 5,000 GPT-3.5 prompts and \$20.11 for 5,000 GPT-4 prompts, which is approximately 20 times the cost of GPT-3.5. The total run-time for the entire experiment was 121.5 minutes, involving the testing of 10,000 prompts.
@@
For the embedding-based methods, obtaining text embeddings via the OpenAI API is extremely time-efficient and more economical compared to prompt-based methods. When using the embedding API, data transmission involves smaller, fixed-size vectors instead of variable-length text, which reduces data overhead and accelerates processing speed. We spent approximately \$0.18 to obtain text embeddings for all 77,245 headlines, while the prompt experiment with only 10,000 prompts cost around \$20.232323At the time of this study (May 2024), GPT-4 (turbo version) prompt costs \$10 per 1M input tokens and \$30 per 1M output tokens. The embeddings used in our study, text-embedding-3-large, only cost \$0.13 per 1M input tokens and there is no output token cost associated with embeddings. Our entire dataset comprises 77,245 headlines, with an average of 19 tokens per headline, resulting in a total embedding cost of approximately \$0.18. Additionally, processing embeddings is much faster (the maximum Tier 1 rate limit—3,000 requests per minute reached) compared to prompt-based methods (around 82 pairs per minute). This efficiency can lead to lower operational costs and easier integration into real-time applications.
@@
For the LoRA fine-tuning method, the GPU requirements depend on factors such as model size, batch size, optimization algorithm, and floating-point precision. In our experiment, full-parameter fine-tuning without LoRA would require over 100 GB of GPU memory. However, LoRA fine-tuning only requires 30 GB of GPU memory. To illustrate the benefit of reduced GPU requirements, we compare two commonly used GPU models for LLM training. Without LoRA, the NVIDIA H100 GPU, which has nearly the largest GPU memory (80 GB), is insufficient for fine-tuning. In contrast, with LoRA, a more affordable NVIDIA V100 GPU with 32 GB of memory is sufficient. The price of an NVIDIA H100 GPU is ten times higher than an NVIDIA V100 GPU as of May 2024. Thus, using LoRA significantly reduces costs, making fine-tuning more accessible to researchers and practitioners. In our experiment, each replicate involved approximately 2 hours of fine-tuning the Llama-3-8b model on an NVIDIA A100 GPU with 80 GB memory, costing about \$3 on the cloud computing provider AutoDL. We also note that after fine-tuning the model, several cost-efficient techniques are available for deploying it for inference, significantly reducing computational costs and hardware requirements. For instance, llama.cpp allows LLM models to be deployed on machines without a GPU, making implementation more affordable. Moreover, model compression techniques such as quantization, sparsification, and knowledge distillation can drastically cut inference computation costs while maintaining, or only slightly affecting, model performance. For a comprehensive survey on efficient LLM deployment and inference techniques, please refer to (zhou2024survey).
@@
In summary, we recommend using either embedding-based methods or fine-tuning methods, as they achieve similar performance (at least in our setting), superior to prompt-based methods, and are cost-effective. These methods offer different levels of transparency: prompt-based methods using OpenAI’s API are entirely black-box; embedding-based methods use OpenAI’s black-box embedding model combined with a user-controlled classification model; while fine-t tuning open-source models provides full transparency. Fine-tuning methods require more significant engineering effort to fine-tune and deploy an LLM than embedding-based methods.
Appendix F LLM-based CTR Prediction Models
We compare two approaches for CTR prediction: an OpenAI-embedding-based model and a LoRA fine-tuned Llama-3-8b model. In summary, we find the LoRA approach has better performance and we finally choose this model in the first step of our LOLA algorithm.
The OpenAI-embedding-based CTR prediction model is similar to the one used for binary classification, as shown in Figure 3, with three key modifications: changing the loss function from log loss to MSE, using the text embedding of a single headline (3072 dimensions) instead of the concatenated embedding from headline pairs, and switching from binary output to continuous CTR prediction.
To fine-tune the Llama-3-8b model for predicting CTR, we use a slightly different model structure compared to the one used for the pairwise comparison task shown in Figure 5. Specifically, as illustrated in Figure A11, we still use the Llama-3-8b model as the base LLM. The input is a single headline, and we apply a linear transformation to the output vector of the final token from the last Transformer layer. This transformation produces a single value, representing the predicted CTR. The model is trained to minimize the MSE between the real CTR and the predicted CTR. All other hyperparameters, such as batch size, number of training epochs, learning rate and its scheduling, and LoRA-related parameters, remain unchanged as in §$\S$3.3.
Figure A11: Model architecture for CTR prediction based on the Llama-3-8b model.
Figure A12 presents the numerical results, showing both the training loss (MSE) and accuracy throughout the training process. Accuracy is computed as follows: for each dataset (training or test), we use the fine-tuned model to predict the CTR for each headline. For each news article, if the headline with the highest predicted CTR matches the headline with the highest actual CTR, we consider this prediction correct. The accuracy is then calculated as the proportion of news articles for which the model’s predictions are correct.
Figure A12: Performance of fine-tuning the Llama-3-8b model for CTR prediction. The accuracy of the OpenAI embedding-based method and the random guess method are also reported. For the random guess method, the winner headline is uniformly and randomly selected for each news article.
As shown in Figure A12, the fine-tuned Llama-3-8b model significantly outperforms the OpenAI embedding-based method in terms of accuracy. Regarding the R-squared values, the embedding-based approach achieves 0.241, whereas the LoRA fine-tuned method achieves a higher R-squared value of 0.315. In contrast, for the pairwise comparison task, the fine-tuned Llama-3-8b model shows only a marginal improvement over the OpenAI embedding-based method. CTR prediction on all headlines is more challenging than pairwise comparison on significant pairs because the CTR output has higher uncertainty compared to a binary label. We believe this increased difficulty necessitates a more complex model, such as Llama-3-8b, which results in a larger performance gap between the fine-tuned Llama-3-8b and the OpenAI embedding-based method.
Additionally, note that the accuracy on the test set is higher than on the training set. In our experiment, many news articles have headlines with very similar CTRs, making it more difficult to distinguish between them, which lowers the accuracy. This result may be due to having more such news articles in the training set and fewer in the test set. It is possible that with a different random split of the training and test sets, the accuracy of the training set could be higher than that of the test set.
@@ G.1 LOLA with Best Arm Identification
In the implementation, we adopt the same data splitting rule as described in §$\S$4.3. Specifically, we used 70% data to fine-tune the Llama-3-8b CTR prediction model (see details in Web Appendix §$\S$ F), 10% data to fine-tune the $n^{\text{aux}}$ parameter, and tested the performance of the LLM-BAI algorithm on the remaining 20% test dataset. To fine-tune the $n^{\text{aux}}$ parameter, we run the LLM-BAI on the algorithm tuning dataset with different $n^{\text{aux}}$ values in $\{20,50,100,200,300,400,500,700,1000\}$, and choose the one with highest success rate in BAI. We finally choose $n^{\text{aux}}=300$ as the best parameter since it has the highest success rate and the number of pulls is not significantly more than others.
@@
We report the performance of LLM-BAI and the standard benchmarks in Figure A13. Note that the success rate for pure LLM here is higher than the accuracy (46.55%) reported in Figure A12, as we use a more lenient definition of success. Our results indicate that our LOLA algorithm outperforms both the pure BAI method and the A/B test method in terms of success rate when using a similar number of pulls per test. Specifically, LOLA improves the success rate by an absolute 2.39% and saves 1.47% of pulls compared to the A/B test. Compared to the plain BAI method, LOLA improves the success rate by an absolute 2.94% and saves 1.19% of pulls.
@@ G.2 LOLA with Thompson Sampling
Figure A14: Average clicks per experiment per period under different time horizon multipliers. Note that the Y-axis captures the average clicks per test per period. For instance, if there is a test with two headlines conducted under $x=100$, then the average click per period in this test is calculated as $(1+2)/100=0.03$. The Y value is simply the average of this number $0.03$ over all tests. This measure scales well with the platform’s total clicks in tests because headlines in different tests with different numbers of headlines take the same weight in this measure.
@@ G.3 LOLA with Contextual Information
The advantages of LLM-LinUCB are its transparency due to the linearity assumption and ease of execution with only one hyperparameter, as well as the theoretical soundness of its optimal regret results. However, there are two issues with the linearity assumption. First, although we have shown that linear functions almost fully leverage text embeddings from OpenAI to achieve high accuracy (there is almost no difference compared to MLP), the embedding itself sacrifices a lot of information, as demonstrated by the significant outperformance of LoRA fine-tuning over embedding-based classification. Second, once concatenated with user contexts, it is unclear whether the linear function is a good approximation of the rewards model (and this requires further numerical study). In addition, LLM-LinUCB cannot accommodate LoRA fine-tuned open-source LLM, which is a complex nonlinear reward model.

### v3

@@ D.3 Implementation Details
Figure A6: The regression pipeline using the Llama model. Headlines are first tokenized and each token is converted into an embedding vector. For illustrative purposes, only two Transformer layers are shown, although the actual model comprises 32 layers for the 8 billion parameter configuration. The figure highlights the use of uni-directional attention in Llama, whereby each token can only incorporate information from preceding tokens. Connections between Transformer blocks represent the attention mechanism, and only a subset of these links are displayed for simplicity. The output from the last token is fed into a regression head, whose output represents the predicted CTR.
During fine-tuning, our objective is to minimize the mean square error on the training set. We implement LoRA fine-tuning using the PEFT library developed by Hugging Face. We set the dimension of the low-rank matrices as $r=4$. Although LoRA can be applied to any subset of weight matrices in LLMs, applying LoRA to query weight matrices $W_{Q}$ and value weight matrices $W_{V}$ is a common practice and typically gives better performance (hu2021lora). Following this practice, in our experiment, we specifically apply LoRA to the $W_{Q}\in\mathbb{R}^{d\times d_{q}}$ and $W_{V}\in\mathbb{R}^{d\times d_{v}}$ matrices within the Transformer blocks of Llama-3-8b, where $d=4096,d_{q}=4096,d_{v}=1024$. To mitigate over-fitting issues, we also apply a relatively large dropout rate of 0.3 to these low-rank matrices. [^19] Furthermore, we enable 16-bit mixed precision training. Although 16-bit training slightly reduces numerical precision, it accelerates operations on GPUs, which enhances computational efficiency significantly. This mode makes the training process faster and consumes less memory compared to the default 32-bit training, especially for large-scale models like Llama-3-8b.
D.4 Performance on Different Train-test Splitting
Recall that we observe the accuracy on the test set being slightly higher than the accuracy on the training set in Figure 4. This phenomenon is not specific to the LoRA fine-tuning method; the same effect is observed with embeddings. It is likely due to the training-test split resulting in unbalanced samples in different sets. To investigate this issue, we re-run experiments on different splits and report the training and test performance to validate this explanation. To save computation costs, instead of rerunning LoRA fine-tuning on different splits, we use embeddings for this sensitivity check, specifically the OpenAI-256E embedding with the Linear CTR prediction model, as used previously in §$\S$3.2. The split used in the paper corresponds to random seed 42. We reported the results using seeds 1, 2, 3, 4, and 5 in Table A3. The accuracy on the training and test sets is similar across these splits.
Seed 1 Seed 2 Seed 3 Seed 4 Seed 5
Training Set Accuracy 39.98% 40.51% 40.64% 40.50% 40.24%
Test Set Accuracy 40.85% 40.85% 39.84% 40.91% 40.33%
Table A3: Training and test set accuracy for different random seeds.
@@ Appendix E Cost Comparison of Pure-LLM-Based Methods
Prompt \$15.87 (gpt-4o with two demonstrations) 5 mins for 3,263 GPT prompts Low
Embedding \$0.29 (text-embedding-3-large) 8 mins for embedding all 12,039 headlines in test set Moderate
@@
Table A4: Summary of cost, run-time, and effort for different methods. Note that the reported run-time for Prompt-based (GPT-3.5 and GPT-4) and embedding-based methods refer to the time needed to obtain responses or embeddings from OpenAI. Note that the prompt and embedding response time can vary based on the API usage tiers, with our results specifically derived from Tier 1. Since we only make prompt inquiries for test data, to maintain a fair comparison, we report the runtime for the embedding method only for headlines in the test data, although we also need to obtain embeddings for headlines in the training data.
For the prompt-based method utilizing the OpenAI API, the cost was \$2.16 for gpt-3.5-turbo-0125 and \$15.87 for gpt-4, on the test dataset with 3,263 prompts when using two demonstrations. With gpt-4 approximately 7.5 times the cost of GPT-3.5. This cost is calculated based on the length (the number of tokens) of the prompt. Runtime for 3,263 prompts is about 5 minutes when using parallelization. There is an inquiry rate/speed limit set by OpenAI that depends on the account tier. Using parallelism could significantly speed up the runtime until the inquiry rate limit is reached.
@@
For embedding-based methods, obtaining text embeddings via the OpenAI API is more economical than using prompt-based methods. At the time of this study (Nov 2024), a gpt-4 prompt costs \$2.5 per 1M input tokens and \$10 per 1M output tokens. The embeddings used in our study, text-embedding-3-large, cost only \$0.13 per 1M input tokens, with no output token cost associated with embeddings. The embedding cost of our entire dataset with 77,245 headlines is \$0.29. The reported monetary cost for gpt-4 is based on our prompt design, which includes not only the headline but also a description of our problem. Additionally, processing embeddings is much faster (the maximum Tier 1 rate limit—3,000 requests per minute reached) compared to prompt-based methods (500 requests per minute). This efficiency can lead to lower operational costs and easier integration into real-time applications.
@@
For the LoRA fine-tuning method, the GPU requirements depend on factors such as model size, batch size, optimization algorithm, and floating-point precision. In our experiment, full-parameter fine-tuning without LoRA would require over 100 GB of GPU memory. However, LoRA fine-tuning only requires 30 GB of GPU memory. To illustrate the benefit of reduced GPU requirements, we compare two commonly used GPU models for LLM training. Without LoRA, the NVIDIA A100 GPU, which has nearly the largest GPU memory (80 GB), is insufficient for fine-tuning. In contrast, with LoRA, a more affordable NVIDIA V100 GPU with 32 GB of memory is sufficient. The price of an NVIDIA A100 GPU is five times higher than an NVIDIA V100 GPU as of May 2024. Thus, using LoRA significantly reduces costs, making fine-tuning more accessible to researchers and practitioners. In our experiment, each replicate involved approximately 2 hours of fine-tuning the Llama-3-8b model on an NVIDIA A100 GPU with 80 GB memory, costing about \$3 on the cloud computing provider AutoDL. We also note that after fine-tuning the model, several cost-efficient techniques are available for deploying it for inference, significantly reducing computational costs and hardware requirements. For instance, llama.cpp allows LLM models to be deployed on machines without a GPU, making implementation more affordable. Moreover, model compression techniques such as quantization, sparsification, and knowledge distillation can drastically cut inference computation costs while maintaining, or only slightly affecting, model performance. For a comprehensive survey on efficient LLM deployment and inference techniques, please refer to (zhou2024survey).
@@
In summary, we recommend using either embedding-based methods or fine-tuning methods, as they achieve similar performance (at least in our setting), superior to prompt-based methods, and are cost-effective. These methods offer different levels of transparency: prompt-based methods using OpenAI’s API are entirely black-box; embedding-based methods use OpenAI’s black-box embedding model combined with a user-controlled CTR prediction model; while fine-tuning open-source models provides full transparency. Fine-tuning methods require more significant engineering effort to fine-tune and deploy an LLM than embedding-based methods.
Appendix F Sensitivity Analysis for Hyperparameter $n^{\text{aux}}$ in LOLA
In this part, we document the sensitivity analysis results for the hyperparameter $n^{\text{aux}}$. We begin by fixing the other hyperparameter $\alpha=0.08$, which has been fine-tuned for our setting, and focus solely on $n^{\text{aux}}$. The sensitivity results are visualized in Figure A7. To investigate its effect on the final average clicks, we select a wide range of values for $n^{\text{aux}}\in\{0,200,1000,1800,3000,5000\}$. Note that this experiment is conducted on the test dataset. When $n^{\text{aux}}=0$, LOLA essentially reduces to the standard UCB, which performs the worst. We observe that LOLA performs best when $n^{\text{aux}}=1000$. This is why we fine-tuned $n^{\text{aux}}$ within the range $\{600,800,1000,1200,1400\}$ around 1000 in our algorithm. This analysis highlights that the performance of LOLA is robust to the choice of $n^{\text{aux}}$, as $n^{\text{aux}}=1800$ still yields results close to those at $n^{\text{aux}}=1000$.
Figure A7: Result of Sensitivity Analysis for Hyperparameter $n^{\text{aux}}$
@@ G.1 LOLA with Best Arm Identification
In the implementation, we adopt the same data splitting rule as described in §$\S$4.3. Specifically, we used 70% data to LoRA fine-tune the Llama-3-8b CTR prediction model, 10% data to fine-tune the $n^{\text{aux}}$ parameter, and tested the performance of the LLM-BAI algorithm on the remaining 20% test dataset. To fine-tune the $n^{\text{aux}}$ parameter, we run the LLM-BAI on the algorithm tuning dataset with different $n^{\text{aux}}$ values in $\{20,50,100,200,300,400,500,700,1000\}$, and choose the one with highest success rate in BAI. We finally choose $n^{\text{aux}}=300$ as the best parameter since it has the highest success rate and the number of pulls is not significantly more than others.
@@
We report the performance of LLM-BAI and the standard benchmarks in Figure A8. Note that the success rate for pure LLM here is higher than the accuracy (46.86%) reported in Figure 4, as we use a more lenient definition of success. Our results indicate that our LOLA algorithm outperforms both the pure BAI method and the A/B test method in terms of success rate when using a similar number of pulls per test. Specifically, LOLA improves the success rate by an absolute 2.39% and saves 1.47% of pulls compared to the A/B test. Compared to the plain BAI method, LOLA improves the success rate by an absolute 2.94% and saves 1.19% of pulls.
@@ G.2 LOLA with Thompson Sampling
Figure A9: Average clicks per experiment per period under different time horizon multipliers. Note that the Y-axis captures the average clicks per test per period. For instance, if there is a test with two headlines receiving 1 and 2 clicks, conducted under $x=100$, then the average click per period in this test is calculated as $(1+2)/100=0.03$. The Y value is simply the average of this number $0.03$ over all tests. This measure scales well with the platform’s total clicks in tests because headlines in different tests with different numbers of headlines take the same weight in this measure.
@@ G.3 LOLA with Contextual Information
The advantages of LLM-LinUCB are its transparency due to the linearity assumption and ease of execution with only one hyperparameter, as well as the theoretical soundness of its optimal regret results. However, there are two issues with the linearity assumption. First, although we have shown that linear functions almost fully leverage text embeddings from OpenAI to achieve high accuracy (there is almost no difference compared to MLP), the embedding itself sacrifices a lot of information, as demonstrated by the significant outperformance of LoRA fine-tuning over the embedding-based approach. Second, once concatenated with user contexts, it is unclear whether the linear function is a good approximation of the rewards model (and this requires further numerical study). In addition, LLM-LinUCB cannot accommodate LoRA fine-tuned open-source LLM, which is a complex nonlinear reward model.